A board 28 feet long is cut into two pieces, and the lengths of the two piecs are in ratio of 2 to 5. Find the lengths of the two pieces.
Answer by prabhjyot(164)

Assume one piece=2x then the other piece =5x
therefore 2x+5x=28 --->x=4
thus
2x=2*4=8ft
and
5x=5*4=20ft